OpenStudy (anonymous):

if log 4 = 0.602 and log 7=0.845 solve log 1/4 and log 49/64

OpenStudy (tkhunny):

\(\log(1/4) = -\log(4)\) \(\log(49/64) = \log(49) - \log(64) = \log\left(7^{2}\right) - \log\left(4^{3}\right) \) Can you take it from there?

OpenStudy (anonymous):

yes Thank you so much!

OpenStudy (tkhunny):

You have to look around, particularly when you are given an odd piece of information. Seriously, why would you need log(7)? There has to be a reason.